OpenStudy (anonymous):

find the solution set. -6m greater than equals to -6

OpenStudy (anonymous):

-6m >= -6... divide by -6 both sides

OpenStudy (anonymous):

\[-6m\geq-6\] \[m\leq \frac{-6}{-6}\] \[m\leq1\]

OpenStudy (anonymous):

gotta switch the inequality because -6 is negative

OpenStudy (anonymous):

answer [-infinity;1]

OpenStudy (anonymous):

or you could write \[-6m\geq -6\] \[-6m+6\geq 0\] \[6\geq 6m\] \[1\geq m\]